Shirts sell for $30 each but if you buy 3 you receive a 20% discount. How much will you pay for 3 shirts.
Answer by rwm(914)

If you get a 20% discount you pay 80%
.8*3*30
4/5*3*30
4*3*6=12*6
72
or
20% of 30=6
3*6=18
3*30=90
90-18=72